# Goldilocks Isolation: High Performance VMs with Edera

### v1

@@ Abstract
Organizations run applications on cloud infrastructure shared between multiple users and organizations. Popular tooling for this shared infrastructure, including Docker and Kubernetes, supports such multi-tenancy through the use of operating system virtualization. With operating system virtualization (known as containerization), multiple applications share the same kernel, reducing the runtime overhead. However, this shared kernel presents a large attack surface and has led to a proliferation of container escape attacks in which a kernel exploit lets an attacker escape the isolation of operating system virtualization to access other applications or the operating system itself. To address this, some systems have proposed a return to hypervisor virtualization for stronger isolation between applications. However, no existing system has achieved both the isolation of hypervisor virtualization and the performance and usability of operating system virtualization.
@@
We present Edera, an optimized type 1 hypervisor that uses paravirtualization to improve the runtime of hypervisor virtualization. We illustrate Edera’s usability and performance through two use cases. First, we create a container runtime compatible with Kubernetes that runs on the Edera hypervisor. This implementation can be used as a drop-in replacement for the Kubernetes runtime and is compatible with all the tooling in the Kubernetes ecosystem. Second, we use Edera to provide driver isolation for hardware drivers, including those for networking, storage, and GPUs. This use of isolation protects the hypervisor and other applications from driver vulnerabilities. We find that Edera has runtime comparable to Docker with .9% slower cpu speeds, an average of 3% faster system call performance, and memory performance 0-7% faster. It achieves this with a 648 millisecond increase in startup time from Docker’s 177.4 milliseconds.
1 Introduction
@@
Cloud computing allows applications to share infrastructure, reducing the cost of deployment. By offloading the running of applications, organizations can use hardware more efficiently, scale resource usage up and down with traffic, and provision their software quickly. In the early days of cloud computing, applications were run on virtual machines [52], which gave each application its own operating system (OS) kernel. Developers could run applications in this virtual machine, with a hypervisor managing the hardware sharing between many virtual machines. Now, it is more common to run multiple applications on the same kernel using operating system virtualization (OS virtualization) to separate the runtimes of each application [36, 51, 50]. This transition is due to both the efficiency of OS virtualization and the improved ease of use from a large ecosystem of tools available to use OS virtualization to run containers.
@@
Containers are an abstraction over applications used by orchestration frameworks like Docker and Kubernetes to facilitate the use of OS virtualization. In this abstraction, the OS kernel runs a container engine that manages the creation and operation of multiple isolated runtimes, known as containers. This is a powerful abstraction that has been used to add efficiency, monitoring, and security to OS virtualization. By 2027, Gartner predicts that 90% of organizations will be running containerized applications in production [12]. Many of these organizations use OS virtualization in multi-tenant environments, with containers from different users running on the same host OS. A company may set up a collection of containers on a host OS to manage multiple copies of their service, with each customer assigned to one of these copies. Although resource sharing enabled by OS virtualization has improved application runtime and resource usage, it comes at the cost of strong isolation between applications.
@@
OS virtualization uses kernel isolation techniques to isolate applications. These techniques, including namespaces [34], cgroups [33], seccomp [35], and capabilities [32], limit an application’s access to kernel resources such as system calls, privileges, and hardware to allow the application to run independently of other containers. However, the nature of the shared kernel leaves a large attack surface for container escapes in which an attacker running an application in a container is able to gain access to the kernel or other containers, “escaping” the isolation of the container runtime. If a cloud service provider gives a container to each customer, a container escape would allow an attacker to access or manipulate other customers’ sensitive user data. These attacks are frequent [1, 24, 43, 15, 16], with five novel container escapes found in 2022 [9, 7, 6, 10, 8].
@@
In order to prevent container escapes, virtualized applications need stronger isolation, without sacrificing the runtime and ease-of-use present in modern virtualization systems. We find that container escape vulnerabilities include a kernel exploit that an attacker uses to circumvent kernel isolation techniques. These attacks can be prevented by moving the shared kernel out of the trusted computing base (TCB). Hypervisor virtualization, which was used for cloud computing before the creation of OS virtualization technologies in the 2000s, provides a kernel for each application, providing stronger application isolation. A hypervisor manages the virtual machines and provides access to hardware resources. By eliminating the shared kernel, an attacker who compromises the kernel will not gain access to the host machine or other applications without also compromising the host machine’s hypervisor. However, any solution that provides strong isolation will not be adopted unless it can match the runtime and ease-of-use of OS virtualization.
@@
Prior work in improving container isolation has made progress toward achieving this combination of strong isolation and ease-of-use. Unikernel approaches [24, 43, 19, 47, 37] combine the application and kernel into a single light-weight workload. However this approach generally requires specifically building applications for use in a unikernel, as well as re-building on every iteration of the application, hindering development speed. Other virtual machine approaches either require specialized hardware [29, 1, 15, 48, 25, 26], such as CPUs that support virtualization extensions, or have high performance costs [27, 1] due to system call indirection. While these approaches have been used for security-critical applications, these limitations have prevented their widespread adoption.
In this work, we observe that paravirtualization [56, 4] can improve the performance of hypervisor virtualization. Paravirtualization reduces the overhead of hypervisors by requiring that guest OSs replace some system calls with hypercalls to the hypervisor, avoiding costly emulation of I/O components. Paravirtualization has been optimized for cloud use cases [23, 5]. We further observe that modern programming language features can reduce the attack surface of a hypervisor by reducing the occurrence of memory bugs in code. Together, these observations allow us to build an optimized hypervisor for running isolated applications.
Based on these observations, we present Edera, a paravirtualized type-1 hypervisor that provides strong isolation between applications, eliminating container escapes. Edera introduces zones that encapsulate each isolated area of the system. Edera uses a microkernel hypervisor running in a hardened root zone to control memory and CPU usage, then runs applications in zones that act as virtual machines, each containing a full Linux kernel. The microkernel, written in MISRA C, presents a small attack surface for applications running within the zones. The runtime, CPU, and memory usage of Edera are similar to that of popular OS virtualized systems, with less than a second of additional startup time.
@@
We test the usability and flexibility of Edera by applying the design to two use cases. First, we develop a container runtime compatible with the Kubernetes specification. Kubernetes is a popular container orchestration platform for managing virtualized applications, and can be used to test Edera on large workloads. Further, Kubernetes compatibility provides interoperability with the full range of tools built by the Kubernetes community for scalability, monitoring, and more. This implementation demonstrates that Edera provides the same ease-of-use as systems with OS virtualization and can be used as a drop-in replacement for popular container runtimes.
Second, we use Edera for driver isolation, moving device drivers out of the hypervisor and into isolated zones. In particular, we develop an isolated Graphics Processing Unit (GPU) driver zone using Edera. The use of cloud GPUs is gaining popularity for the training and running of AI applications. However, bugs in GPU drivers remain common [57, 20, 54] and can be hard to detect [57, 20] leading to memory corruption in the host kernel and container escapes [11]. To reduce the risk of a GPU vulnerability impacting the kernel, we use an Edera zone to isolate GPU drivers from the hypervisor, without requiring changes to GPU drivers or hardware. This isolation prevents any misuse of the GPU from impacting the host kernel. This implementation demonstrates how Edera can be used to isolate device drivers in addition to applications. The same driver isolation model can be applied to drivers for networking, storage, or peripherals.
@@
  • We present Edera, an optimized type-1 microkernel hypervisor with strong isolation between applications.
@@
  • We compare Edera to other virtualization systems in terms of isolation and performance, and find that it has 10.2% slower CPU speed, an average of 8.3% slower memory, and 648 milliseconds of additional startup time compared to using Docker with runc, while providing strong isolation. When compared to systems with strong isolation, Edera outperforms gVisor by all performance metrics and has comparable performance to Kata Containers on a system with virtualization extensions.
@@
  • We implement a drop-in Kubernetes container runtime replacement that gives users the ease-of-use of modern OS virtualization systems. We further present the first system that allows for isolated AI applications on a shared GPU.
2 Background
@@
Figure 1: A summary comparison of different virtualization designs. Note that containers in this diagram use operating system virtualization, sharing a kernel across all applications.
@@
We provide background on the isolation and performance of virtualization techniques. A summary of virtualization designs is shown in fig. 1. We further provide background about Kubernetes and GPU isolation to motivate our use cases.
2.1 Operating system virtualization
@@
Operating system (OS) virtualization allows multiple containers to run on a shared kernel. These containers include an application along with its runtime for independent deployment. Implementations of OS virtualization include Linux containers [36], Solaris zones [51], and Apiary [50]. Containers simplify deployment and scaling of applications by allowing multiple copies of an application to be dynamically added or removed, and a large ecosystem has been built around their use. However, there is no strong isolation between containers running on the same infrastructure.
@@
OS virtualization uses Linux isolation techniques to prevent containers from interfering with each other, including the use of namespaces [34], cgroups [33], seccomp [35], capabilities [32], and SELinux [40]. However, as containers still operate on the same kernel, many attacks have circumvented these isolation techniques to gain access to the kernel or other containers running on the same machine [1, 24, 43, 15, 16, 9, 7, 6, 10, 8]. To prevent these container escape attacks, systems require stronger isolation.
2.2 Hypervisor virtualization
@@
In contrast to OS virtualization, hypervisor (or virtual machine) virtualization uses a separate “guest” OS for each “virtual machine” (VM) with a hypervisor creating and monitoring multiple VMs. This provides a separate kernel for each application, ensuring stronger isolation. In a type 1 (or bare metal) hypervisor, the hypervisor runs directly on the host hardware. Examples include Xen [4] and VMware. Type 2 (or hosted) hypervisors such as VirtualBox instead run as a process on a host OS. Many hypervisors use x86 virtualization, emulating all hardware that the guest OS expects. To reduce the overhead of virtualization, hypervisors with paravirtualization [56, 4] instead require the guest OS to replace system calls with hypercalls to the hypervisor. This hypervisor-aware design avoids costly emulation of I/O hardware components. Some hypervisor virtualization systems, such as Kernel Virtual Machine (KVM), utilize virtualization extensions available on some modern Intel VT or AMD-V CPUs, allowing for more efficient hypervisor virtualization. However such virtualization extensions are not available on all CPUs.
2.3 Isolation Comparison
@@
Unlike hypervisor virtualization, OS virtualization includes the shared kernel in the trusted computing base (TCB). By including the kernel in the TCB, OS virtualization has weaker isolation. This difference can be seen in practice in the abundance of container escape attacks on systems using OS virtualization [9, 7, 6, 10, 8, 11]. These attacks occur despite the use of Linux isolation techniques because the kernel is part of the TCB, and applications are thus vulnerable to flaws in the kernel design. The shared OS and larger TCB increases the attack surface of OS virtualization, allowing privilege escalation, side channel attacks, and information leakage between containers. Despite this, OS virtualization is common in practice because of its speed and the large ecosystem built around container orchestration technologies like Kubernetes.
2.4 Additional Isolation Techniques
@@
Hypervisor virtualization. Several systems improve the efficiency of hypervisor virtualization. Unikernel approaches [24, 43, 19, 47, 37] use a minimized, per-application kernel to reduce the overhead of creating VMs. This approach generally (with the exception of HermiTux [47]) requires re-building images specifically for the unikernel, which can be slow and may be impossible to do for proprietary applications without source code access. A unikernel approach also means that changes to images require re-compilation of the unikernel, hampering development speed.
Hardware virtualization extensions can also be used to improve efficiency of hypervisor virtualization. Kata Containers [29], Firecracker [1], PVM [15], and RunD [22] use hardware virtualization extensions to reduce virtualization overhead. However, the reliance on virtualization extensions means these systems are not able to run on any modern CPU.
@@
OS Virtualization. Other isolation techniques add isolation layers to OS virtualization. These include gVisor [27] which implements user-level kernel isolation, and vKernel [16] which allows for container-specific security rules. However, these systems do not achieve the isolation of hypervisor virtualization as they still contain a shared kernel.
@@
Dedicated hardware. Other approaches to improving container isolation require dedicated hardware. Oxide computers [48] and Metalvisor [25] use dedicated hardware to more securely run containers. Constellation [26] uses a confidential VM for each Kubernetes cluster, with the cluster encrypted in an enclave during runtime.
2.5 Kubernetes

### v2

@@ Abstract.
Organizations run applications on cloud infrastructure shared between multiple users and organizations. Popular tooling for this shared infrastructure, including Docker and Kubernetes, supports such multi-tenancy through the use of operating system virtualization. With operating system virtualization (known as containerization), multiple applications share the same kernel, reducing the runtime overhead and allowing for many applications to run on the same machine. However, this shared kernel presents a large attack surface and has led to a proliferation of container escape attacks in which a kernel exploit lets an attacker escape the isolation of operating system virtualization to access other applications or the operating system itself. To address this, some systems have proposed a return to hypervisor virtualization for stronger isolation between applications. However, no existing system has achieved both the isolation of hypervisor virtualization and the performance and usability of operating system virtualization.
@@
We present Edera, a VMM built for containers that uses paravirtualization, a minimal hypervisor, and dynamic resource management to enable modern cloud systems to use hypervisor virtualization without sacrificing performance. We illustrate Edera’s usability and performance by creating a container runtime compatible with Kubernetes that runs on the Edera hypervisor. This implementation can be used as a drop-in replacement for the Kubernetes runtime and is compatible with all the tooling in the Kubernetes ecosystem. We find that Edera has runtime comparable to Docker with 2.14% slower cpu speeds and faster memory performance. It achieves this with a 740 millisecond increase in startup time from Docker’s 381 milliseconds.
1. Introduction
@@
Cloud computing allows applications to share infrastructure, reducing the cost of deployment. Using cloud resources, organizations can use hardware more efficiently, scale resource usage up and down with traffic, and provision their software quickly. In the early days of cloud computing, applications were run on virtual machines (Barham et al., 2003), which gave each application its own operating system (OS) kernel. Developers could run applications on this virtual machine, with a hypervisor managing the hardware sharing between many virtual machines. Now, it is more common to run multiple applications on the same kernel using operating system virtualization (OS virtualization) to separate the runtimes of each application (maintainers, 2024j; Price and Tucker, 2004; Potter and Nieh, 2010). This transition is due to both the efficiency of OS virtualization and the improved ease of use from a large ecosystem of tools available to use OS virtualization to run containers.
@@
Containers are an abstraction over applications that contain the application and its dependencies. They are used by orchestration frameworks like Kubernetes to facilitate the use of OS virtualization. The OS runs a container engine that manages the creation and operation of multiple containers. This is a powerful abstraction that has been used to add efficiency, monitoring, and security to OS virtualization. By 2027, Gartner predicts that 90% of organizations will be running containerized applications in production (Galea and Carpenter, 2024). Many of these organizations use OS virtualization in multi-tenant environments, with containers from different users or companies running on the same host OS. Although resource sharing enabled by OS virtualization has improved application runtime and resource usage, the shared host OS means it comes at the cost of strong isolation between applications.
@@
OS virtualization uses kernel isolation techniques such as namespaces (maintainers, 2024h), cgroups (maintainers, 2024g), seccomp (maintainers, 2024i), and capabilities (maintainers, 2024f) to limit a container’s access to kernel resources. Over time, containers have added more of these isolation techniques, starting with just namespaces and adding layers to create more boundaries around the container runtime. However, the nature of the shared kernel leaves a large attack surface for container escapes in which an attacker running an application in a container is able to gain access to the kernel or other containers, “escaping” the isolation of the container runtime. These attacks are frequent (Agache et al., 2020; Madhavapeddy et al., 2013; Manco et al., 2017; Huang et al., 2023, 2024; CVE-2022-0847, 2022, 2022, 2022, 2022, 2022).
@@
In order to prevent container escapes, virtualized applications need stronger isolation, without sacrificing the runtime and ease-of-use present in modern virtualization systems. We find that container escape vulnerabilities usually include a kernel exploit that an attacker uses to circumvent kernel isolation techniques. These attacks can be prevented by moving the shared kernel out of the trusted computing base (TCB). A return to hypervisor virtualization provides a kernel for each application, providing stronger isolation. By eliminating the shared kernel, an attacker who compromises the container will not gain access to the host machine or other applications without also compromising the hypervisor. However, any solution that provides strong isolation will not be adopted unless it can match the runtime, resource usage, and ease-of-use of OS virtualization.
@@
Previous work in container isolation has made progress toward achieving this combination of strong isolation and ease-of-use. Unikernel approaches (Madhavapeddy et al., 2013; Manco et al., 2017; Kuenzer et al., 2021; Olivier et al., 2019; maintainers, 2024k) combine the application and kernel into a single light-weight workload. However this approach generally requires re-building applications for use in a unikernel, as well as on every iteration of the application, hindering development speed. Other virtual machine approaches either require specialized hardware (maintainers, 2024c; Agache et al., 2020; Huang et al., 2023; Oxide, 2024; Mainsail, 2024; maintainers, 2024a), such as CPUs with virtualization extensions, or have high performance costs (maintainers, 2024b). While these approaches have been used for security-critical applications, these limitations have prevented their widespread adoption.
In this work, we observe three things about the current state of cloud computing. First, containers have added many layers of segmentation to act more like a partition of computing resources, but OS virtualization does not have access to hardware partitioning. Second, paravirtualization (Whitaker et al., 2001; Barham et al., 2003) allows hypervisor virtualization on existing cloud infrastructure, without the need for hardware virtualization extensions. Paravirtualization utilizes hypercalls on guest OSs, replacing some system calls and avoiding costly emulation of I/O components. Paravirtualization has been optimized for cloud use cases (Liu et al., 2024; Bui et al., 2019) making it ideal for improving the isolation of containers. Third, recent upstream Linux support for paravirtualization enables existing kernels to run with hyprevisor awareness, without the need for full hardware emulation.
Based on these observations, we present the Edera VMM, an orchestrator for a type 1 hypervisor that uses a minimal design, memory safety, and dynamic resource management to provide strong isolation between containerized applications without sacrificing performance or usability. Edera uses the Xen (Barham et al., 2003) microkernel hypervisor to initialized a hardened root zone that controls the memory and CPU usage. The use of Xen’s type 1 hypervisor gives Edera hardware backed partitioning that is not supported by either containers or a type 2 hypervisor. This root hardened zone contains novel VMM tooling that utilizes disagregation and dynamic resource management. Disagregation separates resources into isolated environments where possible, including drivers and system services. Dynamic resource management utilizes insight from the type 1 hypervisor into running workloads to perform optimizations. These optimizations include shared read-only kernel pages, matching the resource needs of containers, and preventing crashes due to usage spikes. Edera introduces zones that encapsulate each isolated area of the system, including drivers and the guest operating systems. These zones support minimal kernel images, such as those from the Open Container Initiative (OCI). While OCI images are not full micro kernels, they provide improved performance over full virtual machines. Both the microkernel design and the use of Rust for memory safety reduce the hypervisor attack surface.
@@
We test Edera’s usability and flexibility by implementing it along with a container runtime compatible with the Kubernetes specification. Kubernetes compatibility provides interoperability with the full range of compatible tools for scalability, monitoring, and more. This implementation demonstrates that Edera provides the same ease-of-use as systems with OS virtualization and can be used as a drop-in replacement for popular container runtimes.
@@
  • We present Edera, an optimized VMM built for containers.
@@
  • We assess the isolation and performance of Edera, and find that it has 2.14% slower CPU speed, faster memory performance, and 740 milliseconds of additional startup time compared to using Docker, while providing strong isolation. Compared to systems with strong isolation, Edera outperforms gVisor in and has comparable performance to Kata Containers and Firecracker with Edera performing better on CPU and memory operations and worse on system calls.
@@
  • We implement a drop-in Kubernetes container runtime replacement that gives users the ease-of-use of modern OS virtualization systems.
2. Background
@@
Figure 1. A summary comparison of virtualization designs. The containers in this diagram use OS virtualization.
@@
We provide background on the isolation and performance of virtualization techniques, summarized in fig. 1. We further provide background about Kubernetes to motivate our use case.
2.1. Operating system virtualization
@@
OS virtualization allows multiple containers to run on a shared kernel. These containers include an application along with its runtime. Examples of OS virtualization are Linux containers (maintainers, 2024j), Solaris zones (Price and Tucker, 2004), and Apiary (Potter and Nieh, 2010). Containers simplify deployment and scaling of applications by allowing multiple copies of an application to be dynamically added or removed, and a large ecosystem has been built around their use. However, there is no strong isolation between containers running on the same infrastructure.
@@
OS virtualization uses Linux isolation techniques to prevent containers from interfering with each other, including the use of namespaces (maintainers, 2024h), cgroups (maintainers, 2024g), seccomp (maintainers, 2024i), capabilities (maintainers, 2024f), and SELinux (maintainers, 2024l). However, as containers still operate on the same kernel, many attacks have circumvented these isolation techniques to gain access to the kernel or other containers running on the same machine (Agache et al., 2020; Madhavapeddy et al., 2013; Manco et al., 2017; Huang et al., 2023, 2024; CVE-2022-0847, 2022, 2022, 2022, 2022, 2022). To prevent these container escape attacks, systems require stronger isolation.
2.2. Hypervisor virtualization
@@
Hypervisor virtualization uses a separate “guest” OS for each virtual machine (VM) with a hypervisor creating and monitoring VMs. This provides a separate kernel for each application, ensuring stronger isolation. In a type 1 (or bare metal) hypervisor, such as Xen (Barham et al., 2003) or VMware, the hypervisor runs directly on the host hardware. Type 2 (or hosted) hypervisors such as VirtualBox instead run as a process on a host OS. Many hypervisors use x86 virtualization, emulating all hardware that the guest OS expects. Paravirtualization (Whitaker et al., 2001; Barham et al., 2003) instead requires the guest OS to make hypercalls to the hypervisor for privileged operations. This hypervisor-aware design avoids costly emulation of I/O hardware components.
@@
Isolation Comparison Unlike hypervisor virtualization, OS virtualization includes the shared kernel in the trusted computing base (TCB). With this larger TCB, OS virtualization has weaker isolation. This difference can be seen in practice in the abundance of container escape attacks on systems using OS virtualization (CVE-2022-0847, 2022, 2022, 2022, 2022, 2022, 2022). These attacks occur despite the use of Linux isolation techniques because the kernel is part of the TCB, and applications are thus vulnerable to flaws in the kernel design. The shared OS and larger TCB increases the attack surface of OS virtualization, allowing privilege escalation, side channel attacks, and information leakage between containers. Despite this, OS virtualization is common in practice because of its speed, cost savings, and the large ecosystem built around container orchestration technologies like Kubernetes.
2.3. Additional Isolation Techniques
@@
Hypervisor virtualization. Several systems improve the efficiency of hypervisor virtualization. Unikernels (Madhavapeddy et al., 2013; Manco et al., 2017; Kuenzer et al., 2021; Olivier et al., 2019; maintainers, 2024k) use a minimized, per-application kernel to reduce the overhead of creating VMs. This approach generally (with the exception of HermiTux (Olivier et al., 2019)) requires re-building images specifically for the unikernel, which can be slow and may be impossible for proprietary applications without source code access. A unikernel also means that changes to images require re-compilation of the unikernel, hampering development speed.
Hardware virtualization extensions can improve hypervisor efficiency. Kernel-based Virtual Machine (KVM) is a Linux kernel feature built on hardware virtualization extensions that allows the kernel to function as a hypervisor (Developers, 2025). Firecracker (Agache et al., 2020) and PVM (Huang et al., 2023) are virtual machine monitors (VMMs) that manage the device model and creation of virtual machines on the KVM hypervisor. However, reliance on virtualization extensions in the underlying KVM means that these systems cannot run in clouds that do not support nested virtualization. There are also performance impacts, especially for startup time that have hindered adoption in all but the most security-sensitive applications. Further, KVM is built on the Linux kernel, so includes this whole kernel in the TCB. Operation forwarding attacks on Kata (maintainers, 2024c) and Firecracker demonstrate vulnerabilities with the use of KVM to forward operations to the host kernel (Xiao et al., 2023).
Other hypervisor virtualization techiques have been introduced for the cloud. Xen-blanket (Williams et al., 2012) adds an intermediate compatibility layer to cloud virtual machines. This provides performant interoperability, but does not directly run containers.
@@
OS Virtualization. Other isolation techniques add isolation layers to OS virtualization. These include gVisor (maintainers, 2024b) which implements user-level kernel isolation, and vKernel (Huang et al., 2024) which allows for container-specific security rules. However, these systems do not achieve the isolation of hypervisor virtualization as they still contain a shared kernel and filesystem, which attackers can use to discover sensitive data (Carta, 2025).
@@
Dedicated hardware. Other approaches such as Oxide computers (Oxide, 2024) and Metalvisor (Mainsail, 2024) improve container isolation with dedicated hardware. Constellation (maintainers, 2024a) and Confidential Containers (maintainers, 2025) use hardware trusted execution environments (TEEs) to isolate Kubernetes clusters.
2.4. Kubernetes